# Scalable and consistent embedding of probability measures into Hilbert spaces via measure quantization

### v1

@@ Appendix B Link between the diagonalization of the covariance operator and the Gram matrix of inner-products
$\Sigma(h)=\frac{1}{N}\sum\limits_{i=1}^{n}f_{i}\langle f_{i},h\rangle_{\mathcal{H}}=\lambda h$ (43)
Dividing by $\lambda$ in (43), one obtains:
$h=\sum\limits_{i=1}^{N}\frac{1}{N\lambda}f_{i}\langle f_{i},h\rangle_{\mathcal{H}}=\sum\limits_{i=1}^{N}a_{i}f_{i},$ (44)
where $a_{i}=\frac{1}{N\lambda}\langle f_{i},h\rangle_{\mathcal{H}}\in\mathbb{R}$. Injecting (44) in (43) yields:
$\sum\limits_{i=1}^{N}f_{i}\langle f_{i},\sum\limits_{j=1}^{N}a_{j}f_{j}\rangle_{\mathcal{H}}=N\lambda\sum\limits_{i=1}^{N}a_{i}f_{i}$ (45)
Taking the inner-product of (45) in with $f_{l}$ yields:
@@
$\Leftrightarrow\sum\limits_{i=1}^{N}\langle f_{l},f_{i}\sum\limits_{j=1}^{N}a_{j}\langle f_{i},f_{j}\rangle_{\mathcal{H}}\rangle_{\mathcal{H}}$ $=N\lambda\sum\limits_{i=1}^{N}\langle f_{l},f_{i}\rangle_{\mathcal{H}}$
@@
If we note $K$ the $N\times N$ Gram matrix of inner-products, that is $G_{ij}=\langle f_{i},f_{j}\rangle_{\mathcal{H}}$, and $a=(a_{1},\cdots,a_{N})^{T}\in\mathbb{R}^{N}$, then we can rewrite the previous inequality with matrix notations:
$K^{2}a=N\lambda Ka$ (46)
and $a$ can be found solving
$Ka=N\lambda a$ (47)
or equivalently by the diagonalization of the Gram matrix $K$. From $a$, one can recover the eigenelement $h$ with (44).
@@ C.1 Synthetic dataset : an explicit formulation for the pairwise inner-products in the Gram matrix
Moreover, the optimal transport map between $\rho$ and $\mu^{(i)}$ is made explicit in (23). Note that it is optimal in the sense of (5) since $x\mapsto\Sigma_{i}^{1/2}x+b_{i}$ is the gradient of a convex function, and Brenier’s theorem [9] allows to conclude. Then
@@
$\langle T_{\rho}^{\mu^{(i)}}-\mathrm{id},T_{\rho}^{\mu^{(j)}}-\mathrm{id}\rangle_{L^{2}(\rho)}=\langle b_{i},b_{j}\rangle+\frac{1}{3d}\langle\Sigma_{i}^{1/2}-I_{2},\Sigma_{j}^{1/2}-I_{2}\rangle_{F}.$ (48)
@@ C.2 Visualization of the flow cytometry dataset
In this section, we show in Figures 5 and 6 the projections of the embedded data (by either $K$-LOT or $K$-KME) on the first components of PCA. We see that the first two components already allow to discriminate the flow cytometry measurements according to their labels.
@@
C.3 Earth image dataset

### v3

@@ Appendix B Link between the diagonalization of the covariance operator and the Gram matrix of inner-products
$\Sigma(h)=\frac{1}{N}\sum\limits_{i=1}^{n}f_{i}\langle f_{i},h\rangle_{\mathcal{H}}=\lambda h$ (B.1)
Dividing by $\lambda$ in (B.1), one obtains:
$h=\sum\limits_{i=1}^{N}\frac{1}{N\lambda}f_{i}\langle f_{i},h\rangle_{\mathcal{H}}=\sum\limits_{i=1}^{N}a_{i}f_{i},$ (B.2)
where $a_{i}=\frac{1}{N\lambda}\langle f_{i},h\rangle_{\mathcal{H}}\in\mathbb{R}$. Injecting (B.2) in (B.1) yields:
$\sum\limits_{i=1}^{N}f_{i}\langle f_{i},\sum\limits_{j=1}^{N}a_{j}f_{j}\rangle_{\mathcal{H}}=N\lambda\sum\limits_{i=1}^{N}a_{i}f_{i}$ (B.3)
Taking the inner-product of (B.3) with $f_{l}$ yields:
@@
$\Leftrightarrow\sum\limits_{i=1}^{N}\langle f_{l},f_{i}\sum\limits_{j=1}^{N}a_{j}\langle f_{i},f_{j}\rangle_{\mathcal{H}}\rangle_{\mathcal{H}}$ $=N\lambda\sum\limits_{i=1}^{N}a_{i}\langle f_{l},f_{i}\rangle_{\mathcal{H}}$
@@
If we note $K$ the $N\times N$ Gram matrix of inner-products, that is $K_{ij}=\langle f_{i},f_{j}\rangle_{\mathcal{H}}$, and $a=(a_{1},\cdots,a_{N})^{T}\in\mathbb{R}^{N}$, then we can rewrite the previous inequality with matrix notations:
$K^{2}a=N\lambda Ka$
and $a$ solves the following equation
$Ka=N\lambda a$
which can recovered by the diagonalization of the Gram matrix $K$. From $a$, one can recover the eigenelement $h$ of the covariance operator $\Sigma$ with (B.2).
@@ C.1 Synthetic dataset : an explicit formulation for the pairwise inner-products in the Gram matrix
Moreover, the optimal transport map between $\rho$ and $\mu^{(i)}$ is made explicit in (5.1). Note that it is optimal in the sense of (2.4) since $x\mapsto\Sigma_{i}^{1/2}x+b_{i}$ is the gradient of a convex function, and Brenier’s theorem [9] allows to conclude. Then
@@
$\langle T_{\rho}^{\mu^{(i)}}-\mathrm{id},T_{\rho}^{\mu^{(j)}}-\mathrm{id}\rangle_{L^{2}(\rho)}=\langle b_{i},b_{j}\rangle+\frac{1}{3d}\langle\Sigma_{i}^{1/2}-I_{2},\Sigma_{j}^{1/2}-I_{2}\rangle_{F}.$
@@ C.2 Visualization of the flow cytometry dataset
In this section, we show in Figures 6 and 7 the projections of the embedded data (by either $K$-LOT or $K$-KME) on the first components of PCA. We see that the first two components already allow to discriminate the flow cytometry measurements according to their labels.
@@
C.3 Performance of the methods for different values of $K$
In this section, we display additional experiments on the flow cytometry and Airbus datasets, completing Tables 1 and 2 for different values of $K$.
Table 3: Flow cytometry dataset. LDA classification accuracies and execution times after 10-component PCA on the methods with $K=16$.
Method Accuracy (Lab) Accuracy (Type) Time (s)
$\overline{K}$-LOT 100 85 23
$\tilde{K}$-LOT 100 81 103
Random subset of size $K$ + LOT 100 77 22
Random subset of size $K$ + KME 100 67 14
$\overline{K}$-KME 100 83 15
$\tilde{K}$-KME 100 69 96
KME with RFF 73 44 4524
$K$-Nys-KME 100 71 12
Table 4: Flow cytometry dataset. LDA classification accuracies and execution times after 10-component PCA on the methods with $K=32$.
Method Accuracy (Lab) Accuracy (Type) Time (s)
$\overline{K}$-LOT 100 94 25
$\tilde{K}$-LOT 100 81 166
Random subset of size $K$ + LOT 100 77 23
Random subset of size $K$ + KME 100 69 32
$\overline{K}$-KME 100 83 34
$\tilde{K}$-KME 100 69 174
KME with RFF 75 44 4701
$K$-Nys-KME 100 73 19
Table 5: Flow cytometry dataset. LDA classification accuracies and execution times after 10-component PCA on the methods with $K=128$.
Method Accuracy (Lab) Accuracy (Type) Time (s)
$\overline{K}$-LOT 100 88 32
$\tilde{K}$-LOT 100 81 555
Random subset of size $K$ + LOT 100 79 28
Random subset of size $K$ + KME 100 73 376
$\overline{K}$-KME 100 77 387
$\tilde{K}$-KME 100 71 909
KME with RFF 92 44 5676
$K$-Nys-KME 100 73 65
Table 6: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on the methods with $K=16$.
Method Accuracy Time (s)
$\overline{K}$-LOT 88 15
$\tilde{K}$-LOT 88 228
Random subset of size $K$ + LOT 67 21
Random subset of size $K$ + KME 69 523
$\overline{K}$-KME 76 219
$\tilde{K}$-KME 68 732
$K$-Nys-KME 65 169
Table 7: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on the methods with $K=32$.
Method Accuracy Time (s)
$\overline{K}$-LOT 89 17
$\tilde{K}$-LOT 89 249
Random subset of size $K$ + LOT 72 28
Random subset of size $K$ + KME 70 2031
$\overline{K}$-KME 67 2792
$\tilde{K}$-KME 67 2247
$K$-Nys-KME 65 172
Table 8: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on the methods with $K=128$.
Method Accuracy Time (s)
$\overline{K}$-LOT 88 32
$\tilde{K}$-LOT 88 390
Random subset of size $K$ + LOT 80 57
Random subset of size $K$ + KME 70 31915
$\overline{K}$-KME 67 31765
$\tilde{K}$-KME 66 32287
$K$-Nys-KME 65 437
C.4 Earth image dataset